The resultant of two vectors having magnitude 10N and 8N cannot be:
(A) 2N (B) 9N (C) 1N (D) 7N
x2g700

Grade:11

2 Answers

Khimraj
6 years ago
Resultant of two vector havingmagnitude A and B is given by
= \sqrt{A^{2} +B^{2}+2ABcos\Theta }
So resultant can max value of A+B when \Theta = 0
andcan have min value of A-B when \Theta = 180.
So resultant of two vector having magnitude 10N and 8N will be in range [2,18].
So correct option is (C) 1N.
Hope it clears. If you like anser then please approve it.
Piyush Upadhyay
6 years ago
The maximum resultant of these two vectors will be 18 N and minimum will be 2 Newton. So according to me option C is not possible.Hope it helps 😊☺